OpenStudy (anonymous):

solve 2/x-2 - 1/x^2+x-6 =x/x+3

OpenStudy (anonymous):

x = -1 OR x = 5

OpenStudy (anonymous):

Common Denominator: ((x-5) (x+1))/(x^2+x-6) = 0 Simplify: (2 x+5)/((x-2) (x+3)) = x/(x+3) Then Solve.

OpenStudy (anonymous):

multiply through by LCD which is x^2+x-6 2 (x+3) - 1 = x(x-2) 2x+6 - 1 = x^2 -2x x^2 -4x -5 = 0 (x-5)(x+1) = 0 x = 5 or -1

OpenStudy (anonymous):

note x^2+x-6 = (x+3)(x-2)

OpenStudy (anonymous):

ok with this gwunder?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

great

OpenStudy (anonymous):

x=-3/2